Job Costs

At the end of April, Prichard Company had completed Jobs 50 and 51. Job 50 is for 200 units, and Job 51 is for 500 units.

The following data relate to these two jobs:

On April 20, raw materials were requisitioned for production as follows: 200 units for Job 50 at $12 per unit and 1,200 units for Job 51 at $12 per unit.

During April, Prichard Company accumulated 500 hours of direct labor costs on Job 50 and 1,100 hours on Job 51. The total direct labor was incurred at a rate of $24 per direct labor hour for Job 50 and $16 per direct labor hour for Job 51.

The predetermined factory overhead rate is $11.00 per direct labor hour.

a. Determine the balance on the job cost sheets for Jobs 50 and 51 at the end of April.

Job 50 $
Job 51 $

b. Determine the cost per unit for Jobs 50 and 51 at the end of April. Round all answers to the nearest whole cent.

Job 50 $
Job 51 $
Answer #1

a) Job Cost sheet

Job 50 Job 51
Direct material 200*12 = 2400 1200*12 = 14400
Direct labor 500*24 = 12000 1100*16 = 17600
Overhead 500*11 = 5500 1100*11 = 12100
Total Job Cost 19900 44100

b) Cost per unit

Job 50 Job 51
Direct material 200*12 = 2400 1200*12 = 14400
Direct labor 500*24 = 12000 1100*16 = 17600
Overhead 500*11 = 5500 1100*11 = 12100
Total Job Cost 19900 44100
Unit 200 500
Cost per unit 99.50 88.20